# Metaplectic operators with quasi-diagonal kernels

### v2

@@ 3.2. The kernel of metaplectic operators
    If $D^{T}(R(B)^{\perp})=\ker(B)$,
    (66) $k(x,y)=\mu_{S}e^{i\pi DB^{+}x_{1}\cdot x_{1}}e^{i\pi DC^{T}x_{2}\cdot x_{2}}e^{i\pi B^{+}Ay_{1}\cdot y_{1}}\delta_{y_{2}=D^{T}x_{2}}(y_{2})e^{-2\pi i(B^{+}x_{1}-C^{T}x_{2})\cdot y_{1}}.$
@@
$\times\qquad\int_{\ker(B)^{\perp}}\int_{D^{T}(R(B)^{\perp})}f(y_{1}+y_{2})\delta_{y_{2}=D^{T}x_{2}}(y_{2})e^{i\pi B^{+}Ay_{1}\cdot y_{1}}e^{-2\pi i(B^{+}x_{1}-C^{T}x_{2})\cdot y_{1}}dy_{1}dy_{2}$
$=\mu_{S}e^{i\pi(DB^{+}x_{1}\cdot x_{1}+DC^{T}x_{2}\cdot x_{2})}$
$\times\qquad\int_{\mathbb{R}^{d}}f(y_{1}+y_{2})\delta_{y_{2}=D^{T}x_{2}}(y_{2})e^{i\pi B^{+}Ay_{1}\cdot y_{1}}e^{-2\pi i(B^{+}x_{1}-C^{T}x_{2})\cdot y_{1}}dy_{1}dy_{2}.$
This concludes the proof.
@@ 3.3.3. The general case
We first prove the assertion under the simplifying assumption $D^{T}(R(B)^{\perp})=\ker(B)$. By writing $u=u_{1}+u_{2}\in R(B)\oplus{R(B)^{\perp}}$ and $v=v_{1}+v_{2}\in\ker(B)^{\perp}\oplus{D^{T}(R(B)^{\perp})}$,
$\tilde{k}(x,y)$ $=k\ast\varphi(x,y)=\int_{\mathbb{R}^{d}}k(u,v)\varphi(x-u)\varphi(y-v)dudv$
$=\mu_{S}\iiiint e^{i\pi(DB^{+}u_{1}\cdot u_{1}+DC^{T}u_{2}\cdot u_{2})}\delta_{v_{2}=D^{T}u_{2}}(v_{2})e^{i\pi B^{+}Av_{1}\cdot v_{1}}e^{-2\pi i(B^{+}u_{1}-C^{T}u_{2})\cdot v_{1}}$
$\qquad\qquad\times\varphi(x_{1}-u_{1})\varphi(x_{2}-u_{2})\varphi(y_{1}-v_{1})\varphi(y_{2}-v_{2})du_{1}du_{2}dv_{1}dv_{2}.$
The integral is understood in the distributional sense and it is taken over the four spaces in the decompositions of $u$ and $v$ above, see also (62) and (63) above. Consequently,
(95) $\begin{split}\tilde{k}(x,y)&=\mu_{S}\iiint e^{i\pi(DB^{+}u_{1}\cdot u_{1}+DC^{T}u_{2}\cdot u_{2})}e^{i\pi B^{+}Av_{1}\cdot v_{1}}e^{-2\pi i(B^{+}u_{1}-C^{T}u_{2})\cdot v_{1}}\\ &\qquad\qquad\times\varphi(x_{1}-u_{1})\varphi(x_{2}-u_{2})\varphi(y_{1}-v_{1})\varphi(y_{2}-D^{T}u_{2})du_{1}du_{2}dv_{1},\end{split}$
where the convergence of the integral is due to the Gaussians. Let $V_{1}:\mathbb{R}^{r}\to R(B)$, $V_{2}:\mathbb{R}^{d-r}\to R(B)^{\perp}$ and $W:\mathbb{R}^{r}\to\ker(B)^{\perp}$ be parametrizations so that $V_{1}^{T}V_{1}=I$, $V_{2}^{T}V_{2}=I$ and $W^{T}W=I$. Then,

### v3

@@ 3.2. The kernel of metaplectic operators
    In all the other cases
    (66) $\begin{split}k(x,y)&=\mu_{S}e^{i\pi(DB^{+}x_{1}\cdot x_{1}+DC^{T}x_{2}\cdot x_{2})}\int_{\ker(B)^{\perp}}\delta_{y=z+D^{T}x_{2}}(y)e^{i\pi B^{+}Az\cdot z}e^{-2\pi iz\cdot(B^{+}x_{1}-C^{T}x_{2})}dz,\end{split}$
    as an oscillatory integral.
@@
$\times\qquad\int_{\ker(B)^{\perp}}\int_{D^{T}(R(B)^{\perp})}f(y_{1}+y_{2})\delta_{y_{2}=D^{T}x_{2}}(y_{2})e^{i\pi B^{+}Ay_{1}\cdot y_{1}}e^{-2\pi i(B^{+}x_{1}-C^{T}x_{2})\cdot y_{1}}dy_{1}dy_{2}.$
By writing
(68) $f(y)$ $=\int_{{\mathbb{R}^{2d}}}f(z)e^{(y-z)\eta}dzd\eta,$
we obtain
(69) $\hat{S}f(x)$ $=\int_{\mathbb{R}^{d}}f(z)k(x,z)dz,$
with $k$ given in (66). This concludes the proof.
@@ 3.3.3. The general case
We first prove the assertion under the simplifying assumption $D^{T}(R(B)^{\perp})=\ker(B)$. By writing $u=u_{1}+u_{2}\in R(B)\oplus{R(B)^{\perp}}$ and $v=v_{1}+v_{2}\in\ker(B)^{\perp}\oplus{D^{T}(R(B)^{\perp})}=\ker(B)^{\perp}\oplus\ker(B)$, it follows by (66) that
$\tilde{k}(x,y)$ $=k\ast\varphi(x,y)=\int_{{\mathbb{R}^{2d}}}k(u,v)\varphi(x-u)\varphi(y-v)dudv$
$=\mu_{S}\int_{\mathbb{R}^{d}}\int_{\ker(B)^{\perp}}e^{i\pi B^{+}u_{1}\cdot u_{1}+DC^{T}u_{2}\cdot u_{2}}e^{i\pi B^{+}Az\cdot z}e^{-2\pi iz(B^{+}u_{1}-C^{T}u_{2})}$
$\qquad\qquad\times\varphi(x-u)\varphi(y_{1}-z)\varphi(y_{2}-D^{T}u_{2})dzdu,$
where the simplifying assumption is used to factorize the Gaussians. Since the dependence of integrand on $u_{1}$ and $u_{1}$ tensorizes, we can write
(97) $\begin{split}\tilde{k}(x,y)&=\mu_{S}\iiint e^{i\pi(DB^{+}u_{1}\cdot u_{1}+DC^{T}u_{2}\cdot u_{2})}e^{i\pi B^{+}Az\cdot z}e^{-2\pi i(B^{+}u_{1}-C^{T}u_{2})\cdot z}\\ &\qquad\qquad\times\varphi(x_{1}-u_{1})\varphi(x_{2}-u_{2})\varphi(y_{1}-z)\varphi(y_{2}-D^{T}u_{2})du_{1}du_{2}dz,\end{split}$
where the convergence of the integral is due to the Gaussians, and the integrals are taken over $R(B)$, $R(B)^{\perp}$ and $\ker(B)^{\perp}$. Let $V_{1}:\mathbb{R}^{r}\to R(B)$, $V_{2}:\mathbb{R}^{d-r}\to R(B)^{\perp}$ and $W:\mathbb{R}^{r}\to\ker(B)^{\perp}$ be parametrizations so that $V_{1}^{T}V_{1}=I$, $V_{2}^{T}V_{2}=I$ and $W^{T}W=I$. Then,